Left Front Bent Gregorian Technical Observing with AGW3: UT 2010 March 04

Observer: DMiller JHill (Tucson)
Telescope Operator: WWack
Instrument Support:
Telescope Support:
SW support: NCushing (Tucson)

Summary

Engineering and LBT work occupied the telescope until 6 UT. We are working at LFBG with AGW3 since the AzCam computer in AGW4 is broken.

Then we spent 3 hours completing the on-sky tests of TCS Build AO2. There were lots of minor issues which may or may not be related to AO2. See the large selection and variety of Issuetraks.

After that we spent the rest of the night semi-productively doing assorted tests. We tested Doug's IDL scripts for sending presets and offsets.

Details

Times in UT

Open

06:32 pointto M5_1311 ACQUIRE

06:34 Wayne tweaked the focus and set IE=-42 CA=-5 to center the star (we didn't check IA.)

06:35 pointto BS9143 ACTIVE gs=0

06:47 dozen or so Act Opt corrections and WFE<200 nm rms. Seeing ~0.9"

06:54 pointto BS9142 ACTIVE gs=9 Failed with error: couldn't request a full CCD readout for GS acquisition. Error is -13 IssueTrak # 2559

06:54 pointto BS9142 ACTIVE gs=9 success. star is 15.7 mag needed for AO2 testing. Guide and collimate for 10 minutes

07:07 offset -180 0 Reacquired the guide star. Ooops, what -180 in DEC rather then RA

07:08 offset 0 -180 Re-acquired the guide star. Didn't move far enough

07:09 offset 0 -90 Could move guide probe to that position. Now in TRACK mode

07:11 offset 0 90 Guide star re-acquired. Guiding and Act Opt.

07:12 offset 0 0 type=ABS Guiding and Act Opt.

07:18 pointto BS9143 gs=0 ACTIVE so we have a bright guide star for MODS tests. Preset failed with the same error as at 06:54 error: couldn't request a full CCD readout for GS acquisition. Error is -13

07:20 Authorize MODS on IIF GUI. Success, but GCS thinks MODS is at LDG, but we are using LFGB.

07:23 pointto BS9143 gs=0 ACTIVE Failed no AGW unit selected

07:25 John edited

07:34 pointto BS9143 gs=0 ACTIVE Failed couldn't find a star in image. Realize that MODS does not have a point and collimation model associated with it. Note, in the acquisition image there was not star, but Sextractor found a noise star along the left side.

07:43 John copied pointing and collimation from Lucifer to MODS files

07:45 Wayne authorized Lucifer then MODS to force reading of new pt and col models

07:45 pointo BS9143 gs=0 ACTIVE Failed GCS died unexpectedly IssueTrak # 2561

07:49 pointto BS9143 gs=0 ACTIVE Failed, no guide star. John realized secondary position was not correct so he copied a collimation lookup table from LUCIFER to MODS

07:59 pointto BS9143 gs=0 ACTIVE Failed with error: couldn't request a full CCD readout for GS acquisition. Error is -13

08:11 Authorized LBTI on both sides. Error with wrong number of lines in primary Mirror config data file. John's looking in to it. Found editing error that he made

08:14 Authorized LBTI on both sides. Successful

08:17 Authorized LINC on both sides. Succussful

08:29 John edited the IRC config.server file so single IRTC images are now placed in /newdata

Testing new IDL routines

09:13 Sent preset, BS9143, GS=0, MODE=ACTIVE, POSANGLE=180. Successful, but start if very out of focus

09:18 preset, BS9143, GS=0, MODE=ACQUIRE, ROTMODE=PARALLACTIC Successful. Wayne tweak focus and set IE=-57 CA=0

09:13 Sent preset, BS9143, GS=0, MODE=ACTIVE, POSANGLE=180. Successful. Guiding and Collimating

WCS orientation tests

These images are to connect Position Angle for WCS to rotator angle. (It may not be a large enough set for that as we got distracted by other things.)

09:40 pointto M92astrom TRACK

/OldRepository/AGW_Data/20100304/

guider0003-5 20 sec PA=0 r' parked

09:49 offset 20 0 RADEC ABS

guider0006-8 20 sec PA=0 r' parked

09:51 setGSsfp left

guider0009-11 20 sec r' PA=0 center

focus 0

09:55 pointto M92astrom ACTIVE gs=4 PA=INDEF=340.72 - couldn't find star

09:58 pointto M92astrom gs=4 ACQUIRE

Adjust IE=-36 CA=19 (not certain this is the right star as we are in M92)

10:10 pointto M92astrom ACTIVE gs=4

Seeing degrades to 2 arcsec and humidity rising.

10:14 pointto M92astrom ACTIVE gs=4 PA=320 Error=-13

10:16 pointto M92astrom ACTIVE gs=4 PA=320 - It seems we have the WFS vignetting problem. Image quality is diverging. No, it seems that GCS wasn't sending WFS values.

Then we find a bug in irc when we try to WFS with IDL - Issue 2562.

10:34 pointto M92astrom ACTIVE gs=4 PA=0 couldn't get full CCD readout

10:40 GCSGUI gives no hint of a warning when star is too dim on WFS. Issue 2563.

Has that been this whole problem. Not only are there no WFS corrections sent, but no collimation updates, so we drift out of focus.

GCSGUI sent WFS integration time in msec when it advertises seconds. Issue 2564.

10:51 offset 10 0 RADEC ABS

11:00 Doug is making offset tests from IDL.

11:08 pointto M92astrom ACTIVE gs=4 PA=0 - motor already in motion
   irc PresetTelescope 0. POSITION ACTIVE left
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = OK
Resmsg[2]  = AGw command status: AGW3: Motor already in motion
Resmsg[3]  = error: setProbePosition: couldn't move probe to position (SFP) (x,y,z): -9.33826, 26.1659, 30.587 mm.
Resmsg[4]  = guidestar list successfully changed but probe couldn't be moved to position.

11:09 home AGW3, but it doesn't work. motor 1 is hung.

11:16 stopAGW, startAGW, home

11:20 pointto M92astrom ACTIVE gs=4 PA=0
   irc PresetTelescope 0. POSITION ACTIVE left
Resmsg[0]  = PresetTelescope result status: Error
Resmsg[1]  = OK
Resmsg[2]  = error: couldn't request a full CCD readout for GS acquisition. Error is -103

11:21 pointto M92astrom ACTIVE gs=4 PA=0 - Error -103 again.

Doug tries command line image.

11:22 Stop and Restart GCSL as it must have been hung from restarting AGW3.

11:23 pointto M92astrom ACTIVE gs=4 PA=0 - succeeded, but problems again with wfs star being too faint.

11:30 pointto BS9169 gs=0 but it is also too faint

11:36 pointto BS9177 gs=41 - no star

11:39 pointto M5_0699 ACQUIRE IE=-52 CA=8

11:44 pointto BS9177 gs=41 ACTIVE

11:51 Doug loses star doing offsets with rotator.

11:53 pointto BS9177 gs=41 ACTIVE PA=INDEF=168.77

Check IA

12:04 pointto M5_0555 ACQUIRE - Adjust IE=-49 CA=+5 at EL=79.7

12:09 pointto M5_0384 moved IA from -266 to -273 (within the errors of not changing)

Check WFS vignetting

12:14 pointto BS9149 gs=0 ACTIVE EL=31.75 ROT=317 - seeing is pretty bad down here at EL=32, but we don't see any problems.

Close

12:30 Close for lack of anything particular more to do. T=-4.1 degC D=-8.8 degC.

-- JohnHill - 04 Mar 2010
Topic revision: r3 - 04 Mar 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback